Phenol, C6H5OH, has a Ka of 1.3×10−10.

1.Write out the Ka reaction for phenol.

2. Calculate Kb for phenol’s conjugate base.

3.Is phenol a stronger or weaker acid than water? explain why.


1
Expert's answer
2021-02-28T06:22:00-0500

1. Ka=[H+]•[C6H5O-]/[C6H5OH]

2. Kb=10-14/Ka=10-14/1.3•10-10= 7.7•10-5

3. From the table data it is known that Ka of water is 1.8•10-15, so Ka of phenol is bigger, and phenol is a stronger acid than water.